Please help me, I don't understand I've tried it and came up with -1874571.4 but I know the answer is 714.5. Could you please show me the steps...
*** Find the partial sum of the geometric series. S9 for 1000-400+160-64+...

Answer by rothauserc(4718)

S9 = a +ar +ar^2 +ar^3 +ar^4 + .... +ar^9, where r is the common ratio
S9 = a(1 - r^9) / (1 - r)
We are given 1000 -400 +160 -64 +
To find r, we divide the second term by the first
r = -400 / 1000 = -2/5
S9 = 1000(1 -(-2/5)^9) / (1 -(-2/5)
S9 = 1000(1 -(-0.000262144)) / 1.4
S9 = 1000 * 1.000262144 / 1.4
S9 = 714.47296